# Cardiocirculatory Computational Models for the Study of Hypertension

### v1

@@ 4.1.1 Single-Parameter Sensitivity Analysis by means of Sobol Indices
Sobol indices [47] are a form of global sensitivity analysis: they rely on the decomposition of the variance of the model’s output into components, which correspond to different input variables and their interactions. In this paper, total Sobol indices $\mathcal{S}^{j,\mathcal{T}}_{k}$ have been employed: they take into consideration both the direct effect of a parameter $p_{k}$ and its interactions with the others. The methodology and analysis of this section are directly derived from [48].
@@
In the parameter sampling process, Sobol sequences are used to ensure a more uniform and space-filling coverage of the hypercube, by minimizing clustering and gaps. This quasi-random structure allows for faster convergence of the sample mean to the expected value compared to standard Monte Carlo sampling, thereby reducing the number of simulations needed to explore the parameter space effectively [49]. Sobol indices are computed using Saltelli’s method [50]: this technique guarantees a high-resolution variance decomposition, providing an accurate information on the contribution of each parameter and its interactions. Though computational effort of Saltelli’s method scales linearly with the number of parameters, it is still feasible, since $(\mathscr{C}_{\text{NC}})$ and $(\mathscr{C}_{\text{C}})$ are computationally cheap, and parallelization limits the time taken for processing. In this paper, $\text{N}_{p}=26$ parameters are analyzed for $(\mathscr{C}_{\text{NC}})$, while, in the case of $(\mathscr{C}_{\text{C}})$, the number of parameters taken in account rises up to 32; the method generates $2\text{N}\left(\text{N}_{p}+1\right)$ samples ($\text{N}=2^{12}$ is a user-defined value), resulting in 221,184 samples for $(\mathscr{C}_{\text{NC}})$ and 270,336 samples for $(\mathscr{C}_{\text{C}})$. The most relevant total Sobol indices $(\geq 0.2)$ for $(\mathscr{C}_{\text{NC}})$, obtained through this procedure and subsequently utilized, are shown in Figure 2.
@@ 4.1.2 Multi-Parameter Sensitivity Analysis
All the mentioned parameters undergo modification by the multiplication with coefficients in such a way that they are constrained within realistic intervals, while all the others remain unchanged. The modeling of hypertension by means of lumped-parameter models is a topic that has not been systematically exploited yet; different studies employ different models, each with different structures and peculiarities, such as in [51, 52, 53]. Since our sensitivity analysis focuses on the hypertensive setting, constructing realistic parameter intervals is not a trivial task; in this paper, inspiration has been taken from [51, 52, 53] and Section 2, in order to obtain coefficients that would be acceptable and useful for carrying out the sensitivity analysis. These coefficients have been chosen to reflect the typical parameter variations observed during the progression of hypertension: resistances and inductances are increased while conductances are decreased, and vice versa; active and passive elastance of the left ventricle are enhanced while the resting volume is reduced, and vice versa. Such modifications try to reproduce the haemodynamic alterations reported in clinical and modeling studies, while keeping the parameters within physiologically plausible ranges. The coefficients are defined as follows: The coefficients are defined as follows:
@@ 4.3.1 Calibration of the 0D model
The calibration procedure that is presented in this section follows the framework described in [31], which, in turn, relies on the method presented in [54]. The process begins by defining a loss function that quantifies the discrepancy between simulation results and target values. The aim of calibration is identifying the parameter configuration $\bar{\mathbf{p}}^{j}\in\Theta=\mathbb{R}^{\text{N}_{p}}$, with $\text{N}_{p}$ being the number of parameters to calibrate, that minimizes the loss function; in order to find the configuration $\bar{\mathbf{p}}^{j}$, parameters are iteratively modified within predefined bounds, untill the loss function is sufficiently small. In [31], the loss function is given by the sum of squared relative errors between model outputs and target data, for the specific individual $j$, as follows:
@@
To minimize $\mathcal{L}^{j}$, parameters are optimized within predefined intervals $\text{I}_{i}$, where $i\in\{1,\dots,\text{N}_{p}\}$. These intervals are set around some initial reference values in such a way that, after the optimization, the changes in parameters remain physically and clinically relevant. The value of the loss function falling below some given small threshold, set to $10^{-3}$, is assumed to indicate a successful calibration. Nevertheless, there are cases where the procedure might be unsuccessful, such as when the loss function’s minimum is located above the threshold, meaning the parameter set is not able to reproduce the reference data within the given constraints. The minimization of $\mathcal{L}^{j}$ is performed in Python using the Quasi-Newton optimization algorithm L-BFGS-B [55], as implemented in the SciPy library [56].
@@ 4.3.2 Calibration of the 3D–0D model
As illustrated in Figure 5, the parameters of $(\mathscr{C}_{\text{NC}})$ are calibrated, using the time-independent outputs of $(\mathscr{C}_{\text{C}})$, and then included in the 3D–0D model. It is noteworthy that the parameters of the 3D–0D model related to the electromechanical component of the left ventricle are never modified: as they are very delicate quantities and closely linked to the specific model itself, it would be unfeasible to try to adapt them to simulate hypertension. For this reason, the same parameters defined in [29] are employed. Finally, a simulation is performed under this comprehensive framework, allowing for the analysis of the model’s behaviour under hypertensive conditions. The results obtained from $(\mathscr{C}_{\text{C}})$ and from the 3D–0D model are then compared to check if the calibration is performed successfully: by effectively aligning the outputs of $(\mathscr{C}_{\text{C}})$ and the 3D–0D model, the latter is now able to give real insight into the complex dynamics of hypertension, achieving this goal with minimal computational effort.
@@ 5.1 Systemic Hypertension
Starting from $p_{\text{LV}}$, in Figures LABEL:fig:sys0d:lv-loop and LABEL:fig:sys0d:plv $(\mathscr{C}_{\text{C}})$ shows a marked increase in peak and end-systolic values, reflecting the increased afterload that forces the ventricle to generate higher pressures. The 3D–0D model confirms this trend but also reproduces a more realistic pressure curve, closer to patient data, with the distinct behaviors of pressure in the different phases more clearly visible, together with an increase in both $p_{\text{LV}}$ and $V_{\text{LV}}$ in Figures LABEL:fig:sys3d:lv-loop and LABEL:fig:sys3d:plv: this leads to a more consistent description of ventricular mechanics in hypertensive conditions [23]. As shown in Table 2, the effect on $\text{LV}_{\text{SW}}$ is evident in both models, with a progressive raise as vascular resistance increases. In the 0D case, this increase represents the ventricle’s attempt to compensate for the increase in afterload, even though it proves insufficient to sustain adequate flow. The 3D–0D model also predicts a higher $\text{LV}_{\text{SW}}$, but with a pressure-volume evolution that better matches clinical expectations, as the loop shows an upward shift in the minimal and maximal pressure, thus providing a more accurate picture of the increase in workload [57]. Considering $Q_{\text{AV}}$ in Figures LABEL:fig:sys0d:qav and LABEL:fig:sys3d:qav, both models show a reduction with disease progression, due to compromised filling and limited ejection, with the 3D–0D model displaying more realistic peak values [58], due to the more accurate simulation of systolic ejection dynamics by the 3D ventricle.
@@
As shown in Table 2, CI shows opposite trends in the two models: in $(\mathscr{C}_{\text{C}})$, it appears to increase slightly, mainly due to the greater pressure generation by the ventricle; however, this result is not consistent with clinical evidence [59]. The 3D–0D model avoids this artefact, showing a more stable CI, which is more in line with the haemodynamic response expected in systemic hypertension [59]. Focusing on $p_{\text{AR}}^{\text{SYS}}$ in Figures LABEL:fig:sys0d:parsys and LABEL:fig:sys3d:parsys, both models capture the increase in systolic and diastolic values. In $(\mathscr{C}_{\text{C}})$, this is reflected in an increase in $\text{SAP}_{\text{max}}$ and $\text{SAP}_{\text{min}}$ (Table 2), indicative of the heavier load imposed on the heart; the 3D–0D model confirms these trends, but with a slightly lower systolic peak. Quantitatively, $\text{SAP}_{\text{max}}$ increases steadily, while $\text{SAP}_{\text{min}}$ increases more moderately before showing a slight decrease in the most severe phase (Table 2). With regard to $Q_{\text{AR}}^{\text{SYS}}$ in Figure LABEL:fig:sys0d:qarsys, $(\mathscr{C}_{\text{C}})$ suggests initial stability followed by a decline, which could lead to a reduction in peripheral perfusion. Furthermore, in Figure LABEL:fig:sys3d:parpul the 3D–0D model shows that $p_{\text{AR}}^{\text{PUL}}$ remains nearly unchanged in mild severity, consistently with clinical evidence indicating that systemic hypertension primarily affects the systemic circulation, with pulmonary pressures increasing only in advanced stages due to backward transmission, i.e. the rise in left heart pressures being passively transmitted to the pulmonary vessels.
@@
Finally, active tension in Figure 7 shows a notable peak increase during isovolumetric contraction and ejection, reflecting the need for higher contractile force to counteract the increased afterload; in addition, spatially resolved active tension peaks in the top portion of the left ventricle at $t=0.26\,\text{s}$ and $t=0.35\,\text{s}$, suggesting regional stress, potentially due to local hypertrophy or contractile heterogeneity [60]. The trace of the Green-Lagrange strain tensor $\mathbf{E}=\dfrac{1}{2}(\mathbf{C}-\mathbf{I})$ in Figure 8 shows only minor changes, with a slight increase in $\left|\operatorname{tr}(\mathbf{E})\right|$ maximal values. This indicates a marginal increase in ventricular dilation and a slightly stronger contraction during the cardiac cycle, consistent with the higher arterial pressure. The overall volumetric deformation remains close to the physiological case.
@@ 5.2 Pulmonary Hypertension
Considering the right ventricular PV loop in Figures LABEL:fig:pul0d:rv-loop and LABEL:fig:pul3d:rv-loop, both end-systolic and end-diastolic pressures increase in $(\mathscr{C}_{\text{C}})$, and the ventricle becomes stiffer. The 3D–0D loop shows a similar overall trend, but the 3D representation allows for a more physiologically accurate description of ejection dynamics and regional mechanics [61].
@@
As shown in Table 3, $\text{RV}_{\text{SW}}$ increases in both models, reflecting the increased energy demand imposed by pulmonary hypertension. While $(\mathscr{C}_{\text{C}})$ quantifies this increase as a compensatory response to maintain flow, the 3D–0D model integrates this increase with the spatial distribution of ventricular contraction, providing a more realistic assessment of workload [62]. $p_{\text{AR}}^{\text{PUL}}$ increases in both models (Figures LABEL:fig:pul0d:parpul and LABEL:fig:pul3d:parpul), demonstrating the afterload imposed on the right ventricle; $Q_{\text{AR}}^{\text{PUL}}$ decreases in 0D due to limited ventricular ejection (Figure LABEL:fig:pul0d:qarpul), whereas in the 3D–0D model, the flow curve is flatter and more prolonged, reflecting a delayed peak and a more physiological right ventricular ejection under pulmonary hypertension. [63] (Figure LABEL:fig:pul3d:qarpul). In Figures LABEL:fig:pul0d:qvensys and LABEL:fig:pul3d:qvensys, $Q_{\text{VEN}}^{\text{SYS}}$ is compromised in the 0D model, potentially causing congestion, while in 3D–0D it increases as a compensatory mechanism to attenuate pressure variations, although this may ultimately fail in cases of further aggravated conditions. Finally, $\max\nabla p_{\text{rAV}}$ rises in both models (Table 3), reflecting reduced filling efficiency and reduced overall cardiac performance. Interventricular interactions may further compromise left ventricular filling, an effect present in 0D and less pronounced in 3D–0D due to more accurate spatial resolution.
As pulmonary hypertension mainly affects right ventricular dynamics, only marginal effects are expected on the left ventricle in the absence of left-sided dysfunction [17, 63]; then, no figures related to 3D left ventricular variables are included.
@@ 5.3 Renovascular Hypertension with Secondary Pulmonary Hypertension
Both $p_{\text{AR}}^{\text{SYS}}$ (Figures LABEL:fig:reno0d:parsys and LABEL:fig:reno3d:parsys) and $Q_{\text{AR}}^{\text{SYS}}$ (Figures LABEL:fig:reno0d:qarsys and LABEL:fig:reno3d:qarsys) increase in both models, as the circulation adapts to preserve systemic perfusion. In $(\mathscr{C}_{\text{C}})$, these increases appear more uniform throughout the entire cardiac cycle, while in the 3D–0D model they are more concentrated around the systolic peak: this difference originates from the more detailed spatial representation of the left ventricle in the 3D–0D framework, which captures the non-linear interplay between ventricular contraction, pressure wave propagation and vascular impedance. As a result, the 3D–0D model provides a sharper and more realistic distribution of pressures and flows, whereas $(\mathscr{C}_{\text{C}})$, due to its lumped nature, tends to smooth out temporal variations and therefore depicts a more averaged behaviour. In Figures LABEL:fig:reno0d:parpul and LABEL:fig:reno3d:parpul, $p_{\text{AR}}^{\text{PUL}}$ also increases due to secondary pulmonary hypertension, while the 3D–0D model highlights how the right ventricle buffers left ventricular preload, maintaining higher volumes and systolic work and capturing ventriculo-vascular interactions more accurately than the 0D simulation. In Figure 11, raised $p_{\text{C}}^{\text{PUL}}$ and $Q_{\text{C}}^{\text{PUL}}$ suggest altered lung microcirculation in the 0D representation, risking fluid buildup and impaired gas exchange, while increased $p_{\text{C}}^{\text{SYS}}$ and $Q_{\text{C}}^{\text{SYS}}$ point to systemic microvascular dysfunction, leading to tissue oedema and impaired nutrient delivery [64].

### v2

@@ 4.1.1 Single-Parameter Sensitivity Analysis by means of Sobol Indices
Sobol indices [sensitivity-analysis08] are a form of global sensitivity analysis: they rely on the decomposition of the variance of the model’s output into components, which correspond to different input variables and their interactions. In this paper, total Sobol indices $\mathcal{S}^{j,\mathcal{T}}_{k}$ have been employed: they take into consideration both the direct effect of a parameter $p_{k}$ and its interactions with the others. The methodology and analysis of this section are directly derived from [sobol-tonini].
@@
In the parameter sampling process, Sobol sequences are used to ensure a more uniform and space-filling coverage of the hypercube, by minimizing clustering and gaps. This quasi-random structure allows for faster convergence of the sample mean to the expected value compared to standard Monte Carlo sampling, thereby reducing the number of simulations needed to explore the parameter space effectively [sobol-sequence:discrepancy]. Sobol indices are computed using Saltelli’s method [saltelli-method]: this technique guarantees a high-resolution variance decomposition, providing an accurate information on the contribution of each parameter and its interactions. Though computational effort of Saltelli’s method scales linearly with the number of parameters, it is still feasible, since $(\mathscr{C}_{\text{NC}})$ and $(\mathscr{C}_{\text{C}})$ are computationally cheap, and parallelization limits the time taken for processing. In this paper, $\text{N}_{p}=26$ parameters are analyzed for $(\mathscr{C}_{\text{NC}})$, while, in the case of $(\mathscr{C}_{\text{C}})$, the number of parameters taken in account rises up to 32; the method generates $2\text{N}\left(\text{N}_{p}+1\right)$ samples ($\text{N}=2^{12}$ is a user-defined value), resulting in 221,184 samples for $(\mathscr{C}_{\text{NC}})$ and 270,336 samples for $(\mathscr{C}_{\text{C}})$. The most relevant total Sobol indices $(\geq 0.2)$ for $(\mathscr{C}_{\text{NC}})$, obtained through this procedure and subsequently utilized, are shown in Figure 2.
@@ 4.1.2 Multi-Parameter Sensitivity Analysis
All the mentioned parameters undergo modification by the multiplication with coefficients in such a way that they are constrained within realistic intervals, while all the others remain unchanged. The modeling of hypertension by means of lumped-parameter models is a topic that has not been systematically exploited yet; different studies employ different models, each with different structures and peculiarities, such as in [coeff-sens-analys1, coeff-sens-analys2, coeff-sens-analys3]. Since our sensitivity analysis focuses on the hypertensive setting, constructing realistic parameter intervals is not a trivial task; in this paper, inspiration has been taken from [coeff-sens-analys1, coeff-sens-analys2, coeff-sens-analys3] and Section 2, in order to obtain coefficients that would be acceptable and useful for carrying out the sensitivity analysis. These coefficients have been chosen to reflect the typical parameter variations observed during the progression of hypertension: resistances and inductances are increased while conductances are decreased, and vice versa; active and passive elastance of the left ventricle are enhanced while the resting volume is reduced, and vice versa. Such modifications try to reproduce the haemodynamic alterations reported in clinical and modeling studies, while keeping the parameters within physiologically plausible ranges. The coefficients are defined as follows: The coefficients are defined as follows:
@@ 4.3.1 Calibration of the 0D model
The calibration procedure that is presented in this section follows the framework described in [lumped:capillary], which, in turn, relies on the method presented in [calibration-model]. The process begins by defining a loss function that quantifies the discrepancy between simulation results and target values. The aim of calibration is identifying the parameter configuration $\bar{\mathbf{p}}^{j}\in\Theta=\mathbb{R}^{\text{N}_{p}}$, with $\text{N}_{p}$ being the number of parameters to calibrate, that minimizes the loss function; in order to find the configuration $\bar{\mathbf{p}}^{j}$, parameters are iteratively modified within predefined bounds, untill the loss function is sufficiently small. In [lumped:capillary], the loss function is given by the sum of squared relative errors between model outputs and target data, for the specific individual $j$, as follows:
@@
To minimize $\mathcal{L}^{j}$, parameters are optimized within predefined intervals $\text{I}_{i}$, where $i\in\{1,\dots,\text{N}_{p}\}$. These intervals are set around some initial reference values in such a way that, after the optimization, the changes in parameters remain physically and clinically relevant. The value of the loss function falling below some given small threshold, set to $10^{-3}$, is assumed to indicate a successful calibration. Nevertheless, there are cases where the procedure might be unsuccessful, such as when the loss function’s minimum is located above the threshold, meaning the parameter set is not able to reproduce the reference data within the given constraints. The minimization of $\mathcal{L}^{j}$ is performed in Python using the Quasi-Newton optimization algorithm L-BFGS-B [bfgs], as implemented in the SciPy library [scipy].
@@ 4.3.2 Calibration of the 3D–0D model
As illustrated in Figure 5, the parameters of $(\mathscr{C}_{\text{NC}})$ are calibrated, using the time-independent outputs of $(\mathscr{C}_{\text{C}})$, and then included in the 3D–0D model. It is noteworthy that the parameters of the 3D–0D model related to the electromechanical component of the left ventricle are never modified: as they are very delicate quantities and closely linked to the specific model itself, it would be unfeasible to try to adapt them to simulate hypertension. For this reason, the same parameters defined in [3d-0d_ventr] are employed. Finally, a simulation is performed under this comprehensive framework, allowing for the analysis of the model’s behaviour under hypertensive conditions. The results obtained from $(\mathscr{C}_{\text{C}})$ and from the 3D–0D model are then compared to check if the calibration is performed successfully: by effectively aligning the outputs of $(\mathscr{C}_{\text{C}})$ and the 3D–0D model, the latter is now able to give real insight into the complex dynamics of hypertension, achieving this goal with minimal computational effort.
@@ 5.1 Systemic Hypertension
Starting from $p_{\text{LV}}$, in Figures LABEL:fig:sys0d:lv-loop and LABEL:fig:sys0d:plv $(\mathscr{C}_{\text{C}})$ shows a marked increase in peak and end-systolic values, reflecting the increased afterload that forces the ventricle to generate higher pressures. The 3D–0D model confirms this trend but also reproduces a more realistic pressure curve, closer to patient data, with the distinct behaviors of pressure in the different phases more clearly visible, together with an increase in both $p_{\text{LV}}$ and $V_{\text{LV}}$ in Figures LABEL:fig:sys3d:lv-loop and LABEL:fig:sys3d:plv: this leads to a more consistent description of ventricular mechanics in hypertensive conditions [vasc20]. As shown in Table 2, the effect on $\text{LV}_{\text{SW}}$ is evident in both models, with a progressive raise as vascular resistance increases. In the 0D case, this increase represents the ventricle’s attempt to compensate for the increase in afterload, even though it proves insufficient to sustain adequate flow. The 3D–0D model also predicts a higher $\text{LV}_{\text{SW}}$, but with a pressure-volume evolution that better matches clinical expectations, as the loop shows an upward shift in the minimal and maximal pressure, thus providing a more accurate picture of the increase in workload [lv-hypertrophy]. Considering $Q_{\text{AV}}$ in Figures LABEL:fig:sys0d:qav and LABEL:fig:sys3d:qav, both models show a reduction with disease progression, due to compromised filling and limited ejection, with the 3D–0D model displaying more realistic peak values [qav-syst], due to the more accurate simulation of systolic ejection dynamics by the 3D ventricle.
@@
As shown in Table 2, CI shows opposite trends in the two models: in $(\mathscr{C}_{\text{C}})$, it appears to increase slightly, mainly due to the greater pressure generation by the ventricle; however, this result is not consistent with clinical evidence [ci-syst]. The 3D–0D model avoids this artefact, showing a more stable CI, which is more in line with the haemodynamic response expected in systemic hypertension [ci-syst]. Focusing on $p_{\text{AR}}^{\text{SYS}}$ in Figures LABEL:fig:sys0d:parsys and LABEL:fig:sys3d:parsys, both models capture the increase in systolic and diastolic values. In $(\mathscr{C}_{\text{C}})$, this is reflected in an increase in $\text{SAP}_{\text{max}}$ and $\text{SAP}_{\text{min}}$ (Table 2), indicative of the heavier load imposed on the heart; the 3D–0D model confirms these trends, but with a slightly lower systolic peak. Quantitatively, $\text{SAP}_{\text{max}}$ increases steadily, while $\text{SAP}_{\text{min}}$ increases more moderately before showing a slight decrease in the most severe phase (Table 2). With regard to $Q_{\text{AR}}^{\text{SYS}}$ in Figure LABEL:fig:sys0d:qarsys, $(\mathscr{C}_{\text{C}})$ suggests initial stability followed by a decline, which could lead to a reduction in peripheral perfusion. Furthermore, in Figure LABEL:fig:sys3d:parpul the 3D–0D model shows that $p_{\text{AR}}^{\text{PUL}}$ remains nearly unchanged in mild severity, consistently with clinical evidence indicating that systemic hypertension primarily affects the systemic circulation, with pulmonary pressures increasing only in advanced stages due to backward transmission, i.e. the rise in left heart pressures being passively transmitted to the pulmonary vessels.
@@
Finally, active tension in Figure 7 shows a notable peak increase during isovolumetric contraction and ejection, reflecting the need for higher contractile force to counteract the increased afterload; in addition, spatially resolved active tension peaks in the top portion of the left ventricle at $t=0.26\,\text{s}$ and $t=0.35\,\text{s}$, suggesting regional stress, potentially due to local hypertrophy or contractile heterogeneity [active-tension-various-cond]. The trace of the Green-Lagrange strain tensor $\mathbf{E}=\dfrac{1}{2}(\mathbf{C}-\mathbf{I})$ in Figure 8 shows only minor changes, with a slight increase in $\left|\operatorname{tr}(\mathbf{E})\right|$ maximal values. This indicates a marginal increase in ventricular dilation and a slightly stronger contraction during the cardiac cycle, consistent with the higher arterial pressure. The overall volumetric deformation remains close to the physiological case.
@@ 5.2 Pulmonary Hypertension
Considering the right ventricular PV loop in Figures LABEL:fig:pul0d:rv-loop and LABEL:fig:pul3d:rv-loop, both end-systolic and end-diastolic pressures increase in $(\mathscr{C}_{\text{C}})$, and the ventricle becomes stiffer. The 3D–0D loop shows a similar overall trend, but the 3D representation allows for a more physiologically accurate description of ejection dynamics and regional mechanics [rv-loop-pul-hyper].
@@
As shown in Table 3, $\text{RV}_{\text{SW}}$ increases in both models, reflecting the increased energy demand imposed by pulmonary hypertension. While $(\mathscr{C}_{\text{C}})$ quantifies this increase as a compensatory response to maintain flow, the 3D–0D model integrates this increase with the spatial distribution of ventricular contraction, providing a more realistic assessment of workload [rv-sw-pul-hyper]. $p_{\text{AR}}^{\text{PUL}}$ increases in both models (Figures LABEL:fig:pul0d:parpul and LABEL:fig:pul3d:parpul), demonstrating the afterload imposed on the right ventricle; $Q_{\text{AR}}^{\text{PUL}}$ decreases in 0D due to limited ventricular ejection (Figure LABEL:fig:pul0d:qarpul), whereas in the 3D–0D model, the flow curve is flatter and more prolonged, reflecting a delayed peak and a more physiological right ventricular ejection under pulmonary hypertension. [qarpul-pul-hyper] (Figure LABEL:fig:pul3d:qarpul). In Figures LABEL:fig:pul0d:qvensys and LABEL:fig:pul3d:qvensys, $Q_{\text{VEN}}^{\text{SYS}}$ is compromised in the 0D model, potentially causing congestion, while in 3D–0D it increases as a compensatory mechanism to attenuate pressure variations, although this may ultimately fail in cases of further aggravated conditions. Finally, $\max\nabla p_{\text{rAV}}$ rises in both models (Table 3), reflecting reduced filling efficiency and reduced overall cardiac performance. Interventricular interactions may further compromise left ventricular filling, an effect present in 0D and less pronounced in 3D–0D due to more accurate spatial resolution.
As pulmonary hypertension mainly affects right ventricular dynamics, only marginal effects are expected on the left ventricle in the absence of left-sided dysfunction [pulhyper:guidelines, qarpul-pul-hyper]; then, no figures related to 3D left ventricular variables are included.
@@ 5.3 Renovascular Hypertension with Secondary Pulmonary Hypertension
Both $p_{\text{AR}}^{\text{SYS}}$ (Figures LABEL:fig:reno0d:parsys and LABEL:fig:reno3d:parsys) and $Q_{\text{AR}}^{\text{SYS}}$ (Figures LABEL:fig:reno0d:qarsys and LABEL:fig:reno3d:qarsys) increase in both models, as the circulation adapts to preserve systemic perfusion. In $(\mathscr{C}_{\text{C}})$, these increases appear more uniform throughout the entire cardiac cycle, while in the 3D–0D model they are more concentrated around the systolic peak: this difference originates from the more detailed spatial representation of the left ventricle in the 3D–0D framework, which captures the non-linear interplay between ventricular contraction, pressure wave propagation and vascular impedance. As a result, the 3D–0D model provides a sharper and more realistic distribution of pressures and flows, whereas $(\mathscr{C}_{\text{C}})$, due to its lumped nature, tends to smooth out temporal variations and therefore depicts a more averaged behaviour. In Figures LABEL:fig:reno0d:parpul and LABEL:fig:reno3d:parpul, $p_{\text{AR}}^{\text{PUL}}$ also increases due to secondary pulmonary hypertension, while the 3D–0D model highlights how the right ventricle buffers left ventricular preload, maintaining higher volumes and systolic work and capturing ventriculo-vascular interactions more accurately than the 0D simulation. In Figure 11, raised $p_{\text{C}}^{\text{PUL}}$ and $Q_{\text{C}}^{\text{PUL}}$ suggest altered lung microcirculation in the 0D representation, risking fluid buildup and impaired gas exchange, while increased $p_{\text{C}}^{\text{SYS}}$ and $Q_{\text{C}}^{\text{SYS}}$ point to systemic microvascular dysfunction, leading to tissue oedema and impaired nutrient delivery [capillary-reno].